OpenStudy (anonymous):

how do you find the area, radius, and apothem with the given information; Complete for a regular hexagon with side 18? i have already found the central angle which is 60 degrees. any help would be awesome!

OpenStudy (anonymous):

 this regular hexagon contains 6 congruent equilateral triangles so to find area just take the area of 1 triangle and multiply by 6: Apothem = 9sqrt3 Radius = 18 notice I wrote all this done without any work... I identified I'm working with a 30-60-90 triangle and because of its properties, I automatically know the other two sides. Area 1 triangle = 1/2*base*Apothem = 1/2*18*9sqrt3. So the area of the hexagon is 6*1/2*18*sqrt3 = 54sqrt3